The risk-free rate is 2.5 percent and the market risk premium is 5 percent. Assume that required returns are based on the CAPM. Your $1 million portfolio consists of $ 208 ,000 invested in a stock that has a beta of 1.0 and the remainder invested in a stock that has a beta of 1.2 . What is the required return on this portfolio?
Answer #1

Portfolio beta=(Amount in A*Beta of A+Amount in B*Beta of B)/Total Amount=(208000*1+(1000000-208000)*1.2)/1000000=1.1584000

Required return=risk free rate+beta*market risk premium=2.5%+1.1584000*5%=8.2920%